OpenStudy (anonymous):

What is the sum of the geometric sequence 8, -16, 32 … if there are 15 terms?

OpenStudy (anonymous):

Here: a = 8 and r = -2 And n= 15 here.. So use: \[\large Sum = \frac{a(1 - r^n)}{(1 - r)}\]

OpenStudy (anonymous):

Are you getting or not @sourpatchwatermelon

OpenStudy (anonymous):

no not really. -2^15 = -32768 i know that. If that's how you're supposed to do it.

OpenStudy (anonymous):

Yes I know that method only..

OpenStudy (anonymous):

\[\large Sum = \frac{8(1 + 32768)}{3}\] Solve this..

OpenStudy (anonymous):

so 87,384?

OpenStudy (anonymous):

@waterineyes did I get it right? or?

OpenStudy (anonymous):

Well Done.. Yes you are right..

OpenStudy (anonymous):

thanks